skylarjh23:

A house appreciates (grows in value) about 6% each year. If the house is now worth $158,000, in about how many years will it be worth $200,000? Use the equation 200,000 = 158,000(1.06)x to solve the problem. Round to the nearest year.

sillybilly123:

\(200,000 = 158,000(1.06)^x\) take logs \( \log \dfrac{200}{158} = x \log 1.06 \)